y=3x-6 a.as the of x increases, the value ofy increase
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y = 3x -6 |slope m = 3/1 y=intercept Pt(0,-6)
Note: as y increases/decreases by 3, x increases/decreases by 1
Pt(0,-6) and according to slope Pt(1,-3) and (2,0) are also on this line
CHECKING our Answer***